OpenStudy (anonymous):

A piano recital is scheduled to be ninety minutes long. The piano teacher will play for ten minutes. During the remaining time, the twenty piano students will get equal time for their performance. How long will each piano student perform? Let m represent the number of minutes each piano student performs.

OpenStudy (anonymous):

will each piano student perform? Let m represent the number of minutes each piano student performs. ninety plus twenty times m equals ten twenty times m plus ten equals ninety ten minus twenty times m equals ninety twenty times m equals one hundred

OpenStudy (anonymous):

m=(90-10)/20 20m=90-10 20m+10=90

OpenStudy (anonymous):

the 2nd phrase...